OpenStudy (anonymous):

Maria put the letter tiles a e io and u in a box.each letter is on two separate tiles in the box . What is the probability of randomly selecting a tile with either an a or an e in the box 1/5 2/5 3/5 4/5 Plz show the work and thank you so mich

OpenStudy (anonymous):

Plz need help

OpenStudy (anonymous):

I'm not very good at probability, but here's my try... 5 letters a e i o u 2 tiles per letter = 5 times 2 = 10 tiles number of a tiles and e tiles = 2 + 2 = 4 chance = 4 / 10